# A solution to Banach conjecture

### v1

@@ Abstract.
In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid. Then we give a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
@@ 3. Proof of the main theorem
Lemma 3.1.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If there exists a $H\in G(n,k)$ for $1\leq k\leq n-1$ such that $K\cap H$ and $K\cap H^{\perp}$ are balls, then $K$ is an ellipsoid.
Proof.
Suppose that $K\cap H$ is a $k$-dimensional ball with a radius $r_{1}$, and $K\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
Then, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid where $K\cap\eta^{\perp}\cap H$ is a (k1)(k-1)-dimensional ball with a radius $r_{1}$, and $K\cap\eta^{\perp}\cap H^{\perp}$ is a (nk)(n-k)-dimensional ball with a radius $r_{2}$.
Therefore, for any $\eta\in\mathbb{S}^{n-1}\cap H$, $K\cap\eta^{\perp}$ is an ellipsoid with (k1)(k-1) axes with lengths $r_{1}$ and (nk)(n-k) axes with lengths $r_{2}$, which implies $K$ is a revolution body with respect to $H^{\perp}$; hence, $K$ is an ellipsoid. ∎
Lemma 3.2.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the maximal set $M(K)$ to be
$M(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\geq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
then, for any linearly independent vectors $\theta_{1},\cdots,\theta_{k}\in M(K)$ with $1\leq k\leq n$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ball with a maximal radius.
Proof.
Case I. If $k=1$, $K\cap\mbox{span}(\theta_{1})$ is a $1$-dimensional ball, as $K$ is origin symmetric.
Case II. If $2\leq k\leq n-1$, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ is a $k$-dimensional ellipsoid with $k$ independent vectors $\theta_{1},\cdots,\theta_{k}$ where $\rho_{K}(\theta_{1})=\cdots=\rho_{K}(\theta_{k})$.
Hence, $K\cap\mbox{span}(\theta_{1},\cdots,\theta_{k})$ has to be a $k$-dimensional ball with a maximal radius.
Case III. If $k=n$, by Case II, for any $1\leq i\leq n$, $K\cap\,\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ is a (n1)(n-1)-dimensional ball. For any $\xi\in\mathbb{S}^{n-1}$, if $\xi^{\perp}$ coincides with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for some $i$, $K\cap\xi^{\perp}$ is a (n1)(n-1)-dimensional ball with a maximal radius. If $\xi^{\perp}$ does not coincide with $\mbox{span}\left(\{\theta_{j}\}_{j\neq i}\right)$ for all $i$, we can choose n1n-1 linearly independent vectors in $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ with $i_{1}\neq i_{2}$; otherwise, $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{1}}\right)$ and $\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\mbox{span}\left(\{\theta_{j}\}_{j\neq i_{2}}\right)$ coincide, which gives $i_{1}=i_{2}$.
Therefore, by Case II, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius. Therefore, for any $\xi\in\mathbb{S}^{n-1}$, $K\cap\xi^{\perp}$ has to be a $k$-dimensional ball with a maximal radius, which implies that $K$ is a ball. ∎
It is clear to derive the following two Corollaries from Lemma 3.2.
Corollary 3.1.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then the maximal set $M(k)$ is a $k$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq k\leq n-1$.
Corollary 3.2.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. If we define the minimal set $m(K)$ to be
$m(K)=\{\theta\in\mathbb{S}^{n-1}:\rho_{K}(\theta)\leq\rho_{K}(\eta)\mbox{ for any }\eta\in\mathbb{S}^{n-1}\},$
then, $m(k)$ is a $l$-dimensional sphere in $\mathbb{S}^{n-1}$ for some $0\leq l\leq n-1$.
Lemma 3.3.
Let $K\subset\mathbb{R}^{n}$ satisfy the given condition in Theorem 3.1. Then either the maximal set $M(k)$ is orthogonal to the minimal set $m(K)$ or $M(K)=m(K)=\mathbb{S}^{n-1}$.
Proof.
If $M(K)\neq m(K)$, then $M(K)\cap m(K)=\emptyset$, and thus
$\mbox{dim}(M(K))+\mbox{dim}(m(K))\leq n-2;$
otherwise, $\mbox{dim}(\mbox{span}(M(K)))+\mbox{dim}(\mbox{span}(m(K)))\geq n+1$, where $\mathbb{S}^{n-1}\cap\mbox{span}(M(K))\cap\mbox{span}(m(K))\neq\emptyset$.
Case I. If $M(K)$ is 0-dimensional and $m(K)$ is (n2)(n-2)-dimensional, we choose $\xi\in M(K)$ and consider any (n3)(n-3)-dimensional subsphere $H$ of $m(K)$.
Then $K\cap\mbox{span}(\xi,H)$ forms an ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $H$; hence $\xi\perp H$, which gives $M(K)\perp m(K)$.
Case II. If $M(K)$ is $k$-dimensional ($k\geq 1$) and $m(K)$ is $l$-dimensional ($l\leq n-3$), then for any $\xi\in M(K)$, $K\cap\,\mbox{span}(\xi,m(K))$ is an $l+2$-dimensional ellipsoid, where the ellipsoid attains its maximum at $\xi$ and its minimum in $m(K)$; hence $\xi\perp m(K)$, which gives $M(K)\perp m(K)$. ∎
Proof of Theorem 3.1.
First by Corollary 3.1 and 3.2, we can set the maximal set $M(K)$ to be $k_{1}$-dimensional and $m(K)$ to be $l_{1}$-dimensional. If $M(K)=m(K)=\mathbb{S}^{n-1}$, $K$ is a ball. If not, by Lemma 3.3, we have $M(K)\perp m(K)$ with $k_{1}+l_{1}\leq n-2$.
If $k_{1}+l_{1}=n-2$, by Lemma 3.1, we have $K$ to be ellipsoid.
If $k_{1}+l_{1}<n-2$, we can define an orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{1},\dots,e_{k_{1}+1}\right)=M(K)$ and $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{n-l_{1}},\dots,e_{n}\right)=m(K)$.
Now we can set $M_{1}:=\max\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{1}=\min\{\rho_{K}(\theta):\theta\in\mathbb{S}^{n-1}\}$, and a linear transformation $T_{1}$ to be
$\left(\begin{array}[]{ccc}\frac{m_{1}}{M_{1}}I_{k_{1}+1}&&\\ &I_{n-k_{1}-l_{1}-2}&\\ &&I_{l_{1}+1}.\end{array}\right)$
Take $K_{2}=T_{1}(K)$. It is clear that $K_{2}$ satisfies the given condition in Theorem 3.1, as $T_{1}$ is a linear transformation. Moreover,
$K_{2}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+1},e_{n-l_{1}},\dots,e_{n}\}$
is a $(k_{1}+l_{1}+2)$-dimensional ball with radius $m_{1}$.
Note that, by Corollary 3.1 and 3.2, we can set the maximal set $M(K_{2})$ to be $k_{2}$-dimensional and $m(K_{2})$ to be $l_{2}$-dimensional.
We claim $\mathbb{S}^{n-1}\cap\mbox{span}(M(K),m(K))\subseteq m(K_{2})$. Hence,
$l_{2}\geq k_{1}+l_{1}+1.$
Then, by Lemma 3.3, either $m(K_{2})=\mathbb{S}^{n-1}$ or $M(K_{2})\perp m(K_{2})$.
If $m(K_{2})=\mathbb{S}^{n-1}$, we have $K_{2}$ to be a ball which gives $K$ being an ellipsoid.
If $M(K_{2})\perp m(K_{2})$, we choose the same orthonomal bases $\{e_{1},\dots,e_{n}\}$ such that $\mathbb{S}^{n-1}\cap\mbox{span}\left(e_{k_{1}+1},\dots,e_{k_{1}+k_{2}+2}\right)=M(K_{2})$.
Now we can set $M_{2}:=\max\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$, $m_{2}=\min\{\rho_{K_{2}}(\theta):\theta\in\mathbb{S}^{n-1}\}$ with $m_{2}=m_{1}$, and a linear transformation $T_{2}$ to be
$\left(\begin{array}[]{cccc}I_{k_{1}+1}&&&\\ &\frac{m_{1}}{M_{2}}I_{k_{2}+1}&&\\ &&I_{n-k_{1}-l_{1}-k_{2}-3}&\\ &&&I_{l_{1}+1}.\end{array}\right)$
Take $K_{3}=T_{2}(K_{2})$. It is clear that $K_{3}$ satisfies the given condition in Theorem 3.1, as $T_{2}$ is a linear transformation. Moreover,
$K_{3}\cap\mbox{span}\{e_{1},\dots,e_{k_{1}+k_{2}+2},e_{n-l_{1}},\dots,e_{n}\}$
is a $(k_{1}+k_{2}+l_{1}+3)$-dimensional ball with radius $m_{1}$.
Repeat the above process, we can finally get a ball after finite linear transformations, which means $K$ is an ellipsoid.
proof of the claim. Since $m(K_{2})\leq m_{1}$, if we assume the opposite of the claim, $m(K_{2})<m_{1}$, which means there exists $\eta\in\mathbb{S}^{n-1}$ such that $\rho_{K_{2}}(\eta)<m_{1}$ and $\eta\notin M(K)\cap m(K)$.
Thus, we choose the preimage of $\eta$ under $T_{1}$ to be $\eta$’. Since $k_{1}+l_{1}<n-2$, we have $K\cap\mbox{span}(M(K),\eta^{\prime})$ is at most $k_{1}+2$ dimensional with $k_{1}+2\leq n-1$ thus an ellipsoid with longest axes in $M(K)$ and another axis $b$. Note that $m_{1}$ is the minimum of $K$, then $b\geq m(K)$.
Now we decompose $\rho_{K}(\eta^{\prime})\eta^{\prime}$ into two parts
$\rho_{K}(\eta^{\prime})\eta^{\prime}=M_{1}\eta_{1}+\tilde{b}\eta_{2},$
where $\eta_{1}\in M(K)$ and $\eta_{2}\in\mathbb{S}^{n-1}\cap M(K)^{\perp}$.
And thus, we consider the elliptic $K\cap\mbox{span}(\eta_{1},\eta_{2})$ with long axis $M_{1}$ and short axis $b$. Then $T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ gives another elliptic with long axis $b\,(\geq m_{1})$ and short axis $m_{1}$, where $\rho_{K_{2}}(\eta)\eta\in T_{1}(K\cap\mbox{span}(\eta_{1},\eta_{2}))$ with $\rho_{K_{2}}(\eta)\geq m_{1}$, contradiction.
@@
First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$. Since $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, for any $x\in K\cap\xi_{0}^{\perp}$, we have
$\|\phi_{\xi}(x)\|_{K}$ $=\inf\{\lambda:\phi_{\xi}(x)\in\lambda K\}$
$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(K\cap\xi^{\perp})\}$
$=\inf\{\lambda:\phi_{\xi}(x)\in\lambda(\phi_{\xi}(K\cap\xi_{0}^{\perp}))\}$
$=\inf\left\{\lambda:\phi_{\xi}(x)\in\phi_{\xi}\left(\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right)\right\}$
$=\inf\left\{\lambda:x\in\frac{\lambda}{|\phi_{\xi}|}(K\cap\xi_{0}^{\perp})\right\}$
$=|\phi_{\xi}|\inf\left\{\lambda:x\in\lambda(K\cap\xi_{0}^{\perp})\right\}$
$=|\phi_{\xi}|\|x\|_{K}.$
Then the continuity of the Minkowski gauge holds
$\left|\|\phi_{\xi}(x)\|_{K}-\|\phi_{\eta}(x)\|_{K}\right|\leq\|x\|_{K}\big||\phi_{\xi}|-|\phi_{\eta}|\big|\leq c\|x\|_{K}\|\xi-\eta\|$
for some constant $c$. The second inequality follows from the continuity of the volume of the sections.
Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows, which gives an origin-symmetric star body
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}.$
@@
Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
@@
For any ellipsoid $\tilde{E}$ in $K$ and $\xi\in\mathbb{S}^{n-1}$, $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$ because $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$. As a result of ellipsoid being the intersection body, by Koldobsky and Zhang’s result (see [9] and [13]), we have $|\tilde{E}|\leq|E|$, which implies $E$ is the John ellipsoid of $K$. ∎
@@
By Theorem 2.2, $K$ contain a John ellipsoid $E$ where for any $\xi\in\mathbb{S}^{n-1}$, $E\cap\xi^{\perp}$ is the John ellipsoid of $K\cap\xi^{\perp}$. We can affine transform $K$ to the John’s position $\tilde{K}$, where $E$ becomes a ball $B$. Since John ellipsoid is linear invariant, $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ to be its John ellipsoid, then $\phi_{\xi}$ has to be congruent transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ is $\phi_{\xi}(B\cap\xi^{\perp})$ which is not $B\cap\xi^{\perp}$.
Then $\tilde{K}\cap\xi^{\perp}$ have some volumes for all $\xi\in\mathbb{S}^{n-1}$, and thus by Minkowski-Funk theorem, $\tilde{K}$ has to be a ball, which implies $K$ is an ellipsoid. ∎

### v2

@@ Abstract.
In this paper, we first prove that the origin-symmetric star body with ellipsoidal sections is an ellipsoid. Then, we provide a complete proof of Banach’s isometric subspace problem in finite dimensions through the John ellipsoid.
Key words and phrases:
Ellipsoid characterization, convex body, central sections, John ellipsoid.
@@ 3. Proof of the main theorem
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [13, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
@@
First, we need to show the existence of the star body $E$. It is sufficient to show the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$.
Now we define a set-valued function $\lambda:\mathbb{S}^{n-1}\to GL(n)$ to be
$\lambda(\xi):=\{\phi_{\xi}\in GL(n):K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})\}.$
And we define the linear symmetry group $G_{K\cap\xi_{0}^{\perp}}$ to be
$G_{K\cap\xi_{0}^{\perp}}:=\{g\in GL(n):g(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\}$
Note that for any $\phi_{\xi},\psi_{\xi}\in\lambda(\xi)$, we have
$\psi_{\xi}^{-1}\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp},$
which gives $\phi_{\xi}\in\psi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}}$; that is
$\lambda(\xi)=\phi_{\xi}\circ G_{K\cap\xi_{0}^{\perp}},$
for some $\phi_{\xi}\in\lambda(\xi)$. And thus, the right action of $G_{K\cap\xi_{0}^{\perp}}$ on $GL(n)$ gives the quotient $GL(n)/G_{K\cap\xi_{0}^{\perp}}$, where the set-valued function $\lambda(x)$ becomes a map $\Phi:\mathbb{S}^{n-1}\to GL(n)\backslash G_{K\cap\xi_{0}^{\perp}}$ with
$K\cap\xi^{\perp}=\Phi(\xi)(K\cap\xi_{0}^{\perp})$
Now we need to show $\Phi$ is continuous. Note that the operator norm of $\phi_{\xi}\in\Phi(\xi)$ is bounded, since
$\|\phi_{\xi}\|\leq\frac{\max_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}{\min_{\theta\in\mathbb{S}^{n-1}}\rho_{K}(\theta)}.$
Then we can find an arbitary convergent sequence $\{\xi_{n}\}_{n=1}^{\infty}$ with $\lim_{n\to\infty}\xi_{n}=\xi$ and pick $\phi_{\xi_{n}}\in\Phi(\xi_{n})$ such that there exists a subsequence $\{\xi_{n_{k}}\}_{k=1}^{\infty}$ with $\phi_{\xi_{n_{k}}}$ converging to $\phi\in GL(n)$ in operator norm. And thus, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),K\cap\xi_{n_{k}}^{\perp})+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $d_{H}(\phi(K\cap\xi_{0}),\phi_{\xi_{n_{k}}}(K\cap\xi_{0}^{\perp}))+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp})$
$\leq$ $c\left|\|\phi\|_{op}-\|\phi_{\xi_{n_{k}}}\|_{op}\right|+d_{H}(K\cap\xi_{n_{k}}^{\perp},K\cap\xi^{\perp}).$
As $k\to\infty$, we have
$d_{H}(\phi(K\cap\xi_{0}),K\cap\xi^{\perp})=0,$
which implies $\phi(K\cap\xi_{0})=K\cap\xi^{\perp}$, that is, $\phi\in\Phi(\xi)$. Therefore, $\Phi(\xi_{n_{k}})\to\Phi(\xi)$ as $k\to\infty$.
This means that for any subsequence $\{\Phi(\xi_{n_{k}})\}_{k=1}^{\infty}$, there exists a sub-subsequence $\{\Phi(\xi_{n_{k_{l}}})\}_{l=1}^{\infty}$ that converges to $\Phi(\xi)$. Consequently, we conclude that $\Phi(\xi_{n})\to\Phi(\xi)$.
Moreover, for any $g\in G_{K\cap\xi_{0}^{\perp}}$, $g(J(K\cap\xi_{0}^{\perp}))=J(g(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, where the John ellipsoid of $K\cap\xi_{0}^{\perp}$ is unique. Hence, $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Therefore, the continuity of $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ with respect to $\xi$ follows from the continuity of $\Phi(\xi)$.
Now, we need to show that for any $\xi,\eta\in\mathbb{S}^{n-1}$,
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)$
holds for any $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}\cap\eta^{\perp}$. It is sufficient to show for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
Here we choose $\phi_{\zeta}\in\Phi(\zeta)$ with
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
We partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $\xi$ such that, for every $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$,
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $B$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$. Moreover, the closedness of $A_{1}$ follows the closedness of $G_{K\cap\xi_{0}^{\perp}}$.
For $\xi\in A$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
where
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=J(K\cap\xi_{0}^{\perp})\cap g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
which implies
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(g(J(K\cap\xi_{0}^{\perp}))\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
where $g(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp})\cap\phi_{\zeta}^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}))=g\circ g^{-1}(J(K\cap\xi_{0}^{\perp})\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
that is,
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))\cap\xi_{0}^{\perp}=J(K\cap\xi_{0}^{\perp})\cap\zeta^{\perp}.$
Therefore, we obtain
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$.
For $\xi\in A_{2}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
where
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\xi_{0}^{\perp}).$
Therefore, using the continuity of $\Phi$ and the fact that $\phi^{-1}_{\zeta}\neq g$, for any
$\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta}}{\|(1-t)\xi_{0}+t\eta_{\zeta}\|}\quad\text{with }t\in[0,1],\ \eta_{\zeta}\perp\phi^{-1}_{\zeta}(\xi_{0}^{\perp}),\ \text{and }\eta_{\zeta}\perp\xi_{0},$
there exists a linear map $\phi_{\eta}$ satisfying
$\phi_{\eta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$
and $\phi_{\eta}(\xi^{\perp})=\xi_{0}^{\perp}$. Therefore, we obtain
$\phi^{-1}_{\eta}(K\cap\xi_{0}^{\perp}\cap\eta^{\perp})=K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\eta}(\xi_{0}^{\perp}),$
and thus, $\eta\in B$. Since $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ is compact, we can choose $\zeta_{0}$ such that $\|\eta_{\zeta_{0}},\xi_{0}\|$ reaches the maximum; then $\eta=\frac{(1-t)\xi_{0}+t\eta_{\zeta_{0}}}{\|(1-t)\xi_{0}+t\eta_{\zeta_{0}}\|}$ is in $A_{2}$.
Therefore, for any open set $O\subseteq A_{2}$ and any $\xi,\eta\in O$, there exists a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$. If this were not the case, then by the continuity of $\Phi$ we could find some $\zeta\in O$ such that
$\phi_{\alpha}\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)=g\bigl(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}\bigr)$
for every $\alpha\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi_{0},\zeta)$, where $g\in G_{K\cap\xi_{0}^{\perp}}$. Fixing $\xi\in O$, consider the set
$\{\eta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:K\cap\xi_{0}^{\perp}\cap\eta^{\perp}=\phi_{\eta}(K\cap\xi_{0}^{\perp}\cap\xi^{\perp}),\,\phi\in GL(n)\},$
which is closed. Consequently, for all $\xi,\eta\in\bar{O}$ there is a linear isomorphism between $K\cap\xi_{0}^{\perp}\cap\xi^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\eta^{\perp}$. Then $\bar{O}\subseteq A_{2}$, which shows that $A_{2}$ is closed.
By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $A_{1}$ or $A_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$. If $A_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, the origin-symmetric star body
$E:=\{\phi_{\xi}(J(K\cap\xi_{0}^{\perp})):\xi\in\mathbb{S}^{n-1}\}$
@@
exsits. Then, by Theorem 3.1, $E$ is an ellipsoid in $K$ with $E\cap\xi^{\perp}$ being the John ellipsoid of $K\cap\xi^{\perp}$ for any $\xi\in\mathbb{S}^{n-1}$.
If $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, then any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are mutually linearly dependent; then, using the linear dependence of all (n1)(n-1)-dimensional sections, we deduce that all (n2)(n-2)-dimensional sections of $K$ are likewise linearly dependent; therefore, by Gromov’s theorem [7], $K$ must be an ellipsoid. And hence, $E=K$ is an ellipsoid.
@@
For any ellipsoid $\tilde{E}$ contained in $K$ and any $\xi\in\mathbb{S}^{n-1}$, we have $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, since $E\cap\xi^{\perp}$ serves as the John ellipsoid of $K\cap\xi^{\perp}$. As a result of the ellipsoid being an intersection body, Koldobsky and Zhang’s result (see [9] and [14]) yields $|\tilde{E}|\leq|E|$, and hence $E$ must be the John ellipsoid of $K$. ∎
Remark 3.1.
The proof of the continuity of $\phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ can also be found in [3, Lemma 1.5].
@@
By Theorem 3.2, the body $K$ contains a John ellipsoid $E$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the section $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. By applying an affine transformation, we can move $K$ into John’s position, denoted by $\tilde{K}$, in which $E$ is transformed into a Euclidean ball $B$. Since the John ellipsoid is invariant under linear transformation, the set $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ as its John ellipsoid. Consequently, $\phi_{\xi}$ must be a congruence transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ would be $\phi_{\xi}(B\cap\xi^{\perp})$, which differs from $B\cap\xi^{\perp}$.
Hence, $\tilde{K}\cap\xi^{\perp}$ has the same volume for every $\xi\in\mathbb{S}^{n-1}$, and therefore, by the Minkowski–Funk theorem, $\tilde{K}$ must be a ball, which in turn implies that $K$ is an ellipsoid. ∎
4. Acknowledgement
The crucial idea emerged when the author visited the Tianyuan Mathematical Research Center in Kunming.